OpenStudy (anonymous):

Divide the following polynomial using synthetic division (x^3 + 6x^2 + 3x + 1 ) Ă· (x - 2)

ganeshie8 (ganeshie8):

\( \begin{array}{} 2~|&1&6&3&1\\ &&&\\ \hline &&&&\\ &1&&& \end{array} \)

ganeshie8 (ganeshie8):

multiply 1*2 and put it in next column

ganeshie8 (ganeshie8):

\( \begin{array}{} 2~|&1&6&3&1\\ &&2&\\ \hline &&&&\\ &1&&& \end{array} \)

ganeshie8 (ganeshie8):

add 6 and 2, put it down in last row

ganeshie8 (ganeshie8):

\( \begin{array}{} 2~|&1&6&3&1\\ &&2&\\ \hline &&&&\\ &1&8&& \end{array} \)

ganeshie8 (ganeshie8):

multiply 2*8, put it in next column

ganeshie8 (ganeshie8):

\( \begin{array}{} 2~|&1&6&3&1\\ &&2&16&\\ \hline &&&&\\ &1&8&& \end{array} \)

ganeshie8 (ganeshie8):

add 16 and 3, put it down in last row

ganeshie8 (ganeshie8):

\( \begin{array}{} 2~|&1&6&3&1\\ &&2&16&\\ \hline &&&&\\ &1&8&19& \end{array} \)

ganeshie8 (ganeshie8):

can u continue further ?

OpenStudy (anonymous):

shouldnt it be written something like x^2 + 5x etc...

ganeshie8 (ganeshie8):

yes, finish the synthetic division first

ganeshie8 (ganeshie8):

multiply 2*19, and put it in next column

ganeshie8 (ganeshie8):

\( \begin{array}{} 2~|&1&6&3&1\\ &&2&16&38\\ \hline &&&&\\ &1&8&19& \end{array} \)

ganeshie8 (ganeshie8):

add 38 and 1 and put it in the last row

ganeshie8 (ganeshie8):

\( \begin{array}{} 2~|&1&6&3&1\\ &&2&16&38\\ \hline &&&&\\ &1&8&19& | 39 \\ \end{array} \)

ganeshie8 (ganeshie8):

so, 39 is the remainder

ganeshie8 (ganeshie8):

quotient can be written using the last row, leaving the remainder part

ganeshie8 (ganeshie8):

in the last row u have : \(1~~8~~19\)

ganeshie8 (ganeshie8):

so the quotient is : \(x^2 + 8x + 19\)

ganeshie8 (ganeshie8):

so, \((x^3 + 6x^2 + 3x + 1 ) Ă· (x - 2) ~= ~ x^2 + 8x + 19 + \frac{39}{x-2}\)

ganeshie8 (ganeshie8):

let me knw if smthng doesnt make sense...